lea is 6 years older than sue,but jon is 5 years older than lea, and their ages equal 41, how old is sue?
Answer by stanbon(75887)

lea is 6 years older than sue,but jon is 5 years older than lea, and their ages equal 41, how old is sue?
----------------
Equations:
L = S+6
J = L+5
L + S + J = 41
-------------------
L + (L-6) + (L+5) = 41
---
3L -1 = 41
3L = 42
--------------
L = 14
S = L-6 = 8
J = L+5 = 19
---------------------
Cheers,
Stan H.
====================